# An Investigation of FP8 Across Accelerators for LLM Inference

### v1

@@ 1 Introduction
To fully address the challenges of LLM inference, we set out to quantify the various constraints involved, including those that arise with the inclusion of FP8 quantization. To the best of our knowledge, we are the first to provide quantitative measurements associated with various FP8 quantization schemes.
@@ 4.1 Dynamic vs static scaling
Following on the findings of Shen et al. (2024), which demonstrated that the E4M3 format achieves superior accuracy on language tasks compared to E5M2, we extend this analysis to instruction-tuned models in Table 3. To determine the optimal format for inference, we conducted a comparative evaluation of E4M3 and E5M2, measuring MMLU accuracy using LM Evaluation Harness (v0.4.7, (Gao et al., 2024)). Our experimental results consistently indicate that E4M3 outperforms E5M2 across all evaluated scenarios. Furthermore, as shown in Table 8, the GEMM throughputs using E4M3 and E5M2 are comparable, making E4M3 the preferred choice for inference even considering the smaller representational capacity of E4M3 on Gaudi 2.
@@ 5.1 LLM inference phases
Generative LLM inference consists of a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024). We include a diagram highlighting the differences between the two phases in Figure 2.
In the prefill phase, the LLM generates the first token based on the provided input prompt. The key-value pairs are stored in the KV cache for reuse during the generation of subsequent tokens. The prefill phase processes the entire input in parallel, utilizing matrix-matrix operations that are typically compute-bound. This parallelism allows for effective hardware utilization, enabling efficient input processing.
After the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens. The decode phase involves GEMV operations, inherently limiting hardware utilization compared to the compute-bound GEMMs in the prefill phase.

### v2

@@ 1 Introduction
To comprehensively address the challenges of LLM inference, we aim to quantify the various constraints that impact performance, including those introduced by FP8 quantization. To the best of our knowledge, we are the first to provide quantitative measurements associated with various FP8 quantization schemes.
@@ 4.1 Dynamic vs static scaling
4.2 E4M3 vs. E5M2
Following on the findings of Shen et al. (2024), which demonstrated that the E4M3 format achieves superior accuracy on language tasks compared to E5M2, we extend this analysis to instruction-tuned models in Table 3.
To determine the optimal format for inference, we conducted a comparative evaluation of E4M3 and E5M2, measuring MMLU accuracy using LM Evaluation Harness (v0.4.7, (Gao et al., 2024)). Our experimental results consistently indicate that E4M3 outperforms E5M2 across all evaluated scenarios. Furthermore, as shown in Table 8, the GEMM throughputs using E4M3 and E5M2 are comparable, making E4M3 the preferred choice for inference even considering the smaller representational capacity of E4M3 on Gaudi 2.
@@ 5.1 LLM inference phases
Generative LLM inference comprises two distinct phases: a compute-bound prefill phase and a memory-bound decode phase (Patel et al., 2024; Zhong et al., 2024). Figure 2 illustrates the key differences between these phases.
During the prefill phase, the LLM generates the first token based on the provided input prompt. Throughout this process, key-value (KV) pairs are stored in the KV cache for reuse in subsequent token generations. The prefill phase processes the entire input in parallel, utilizing matrix-matrix operations that are typically compute-bound. This high degree of parallelism allows for effective hardware utilization, enabling efficient input processing.
Once the first token is generated, the process transitions to the decode phase, where output tokens are generated sequentially in an autoregressive manner based on the previous tokens. Unlike the prefill phase, the decode phase is constrained by memory bandwidth, as it involves GEMV operations, which inherently limit hardware utilization compared to the more parallelizable GEMMs in the prefill phase.